A 0.600-π‘˜π‘” particle has a speed of 2.00 π‘š/𝑠 at point 𝐴 and kinetic energy of 7.50 𝐽 at point 𝐡. What is a) its kinetic energy at 𝐴? b) its speed at 𝐡? c) the total work done on the particle as it moves from 𝐴 to 𝐡?


1
Expert's answer
2021-10-26T09:43:41-0400

a)

"KE_A=\dfrac{1}{2}mv_A^2,""KE_A=\dfrac{1}{2}\times0.6\ kg\times(2.0\ \dfrac{m}{s})^2=1.2\ J."

b)

"KE_B=\dfrac{1}{2}mv_B^2,""v_B=\sqrt{\dfrac{2KE_B}{m}}=\sqrt{\dfrac{2\times7.5\ J}{0.6\ kg}}=5\ \dfrac{m}{s}."


c)

"W_{tot}=\Delta KE=KE_b-KE_a,""W_{tot}=7.5\ J-1.2\ J=6.3\ J."